OpenStudy (anonymous):

Find a Power Series Representation for the function and determine the interval of convergence.

OpenStudy (anonymous):

\[f(x)= \frac{ 1+x }{ 1-x }\]

OpenStudy (amistre64):

youll want to start by calculating a few of the derivatives of it

OpenStudy (anonymous):

Low de High - High De Low???

OpenStudy (amistre64):

sounds about right :) or you can up the bottom and just run the product rule

OpenStudy (anonymous):

but then i will still get a fraction... how will i be able to express as a power series

OpenStudy (amistre64):

the derivatives act as the constant terms of all the power terms of the polynomial; if we can spot a pattern we can develop a rule for it

OpenStudy (anonymous):

ok. so i should plug in values?

OpenStudy (amistre64):

do you have the derivatives yet?

OpenStudy (amistre64):

once you have the derivatives, determine a suitable value for x that will be able to turn the derivatives into constants

OpenStudy (anonymous):

For the derivative i got 2/ (1-x^2)

OpenStudy (amistre64):

youll want a few of them, like for or five if you can to develop a pattern

OpenStudy (anonymous):

ohhhhhhh now i see what youre doing. Thanks, i know how to do the problem now. Thanks!

OpenStudy (amistre64):

cool, good luck with the manual labor parts :)

OpenStudy (anonymous):

i would divide first

OpenStudy (anonymous):

that is, start with \[-1+\frac{2}{1-x}\] then the power series for \(\frac{1}{1-x}\) is well known and you can use that one

OpenStudy (anonymous):

you get the answer pretty much right away without using a derivative at all \[1+2x+2x^2+2x^3+...\]

OpenStudy (amistre64):

well yeah, when you know a few tricks :)

OpenStudy (anonymous):

i like tricks